A 5.00 L of a gas is collected at 100. K and then allowed to

expand to 20.0 L. What must the new temperature be in order

to maintain the same pressure (as required by Charles' Law)?


1
Expert's answer
2022-01-25T05:38:01-0500

"\frac{V_1}{T_1}=
\frac{V_2}{T_2}"


"\frac{5.00}{100}=
\frac{20.0}{T_2}"


"T_2=\frac{100×20.0}{5.00}=400K"